An electric immersion heater has a power rating of 1800 W. If the heater is placed in a 1.2 kg of water at 25 °C, how many minutes will it take to bring the water to a boiling temperature? (Assume that there is no heat loss except to the water itself. cwater=4186 J/kg.°C)
Answer #1

From the given question,

power rating of immersion heater(P)=1800w

mass of water(m)=1.2 kg

initial temp of water(T1)=25 deg cel.

final temp of water(T2)=100 deg cel.

Change in temperature(\DeltaT)=100-25=75 deg cel

time=t (in sec)

specific heat of water(c)=4186 J/kgoC

electric energy of heater= heat energy

Pt= mc\DeltaT

1800 x t= 1.2 x 4186 x 75

t=209.3 sec

=3.5 min (approx)

It will take 3.5 minutes to bring water to boiling temperature.